# Fast and accurate analytical formulas for light propagation in general static, spherically symmetric spacetimes

### v1

@@ III.2.1 Einstein-Maxwell-dilaton-axion metric
Figure 4 illustrates the performance of the analytical approximations given by Eqs. (23) and (24) for the Einstein–Maxwell–dilaton–axion metric, considering three representative configurations corresponding to different combinations of the parameters $(b_{0},\mu)$ and the associated $(r_{H},r_{c})$ values. The plots display the relation between $\cos\psi$ and $\cos\alpha$ for several emission radii $R$, together with the relative percentage deviation of $\alpha$ in the lower panels. Across all parameter choices, the full analytical approximation given by Eq. (23) (Approx. T) shows excellent agreement with the exact numerical integration. The discrepancy remains within $|\delta\alpha|/\alpha<3\%$ over the entire domain of $\psi$, even for the smallest radii considered—namely, $R=1.1r_{H}$ and $R=1.01r_{c}$—where strong deflection effects dominate. This shows that the approximation maintains high accuracy even in the deep strong-field regime near the event horizon. The linearized expression derived from Eq. (24) (Approx. L) also performs well for moderate deflections ($\cos\psi\gtrsim 0$), particularly at larger radii ($R\geq 1.5r_{c}$), but gradually loses precision as $\psi\to\pi$ and $\cos\psi\to-1$. This degradation is more evident in the rightmost panel, where the dilaton parameter dominates ($b_{0}>\mu$), leading to a more compact horizon and stronger curvature gradients. In contrast, in the leftmost configuration ($b_{0}<\mu$), both approximations remain nearly indistinguishable from the numerical result except for very large bending angles. As in the previous cases, the lower panels quantify these trends. For all three configurations, Approx. T maintains a smooth and nearly symmetric error distribution around $\cos\psi=0$, indicating robustness under parameter variations. Approx. L, while showing larger residuals near $\cos\psi\approx-1$, still provides a reasonable description in the weak-to-intermediate deflection regime, making it suitable for analytical estimates where computational simplicity is preferred. Overall, these results show that the generalized analytical formulas accurately reproduces light propagation in the EMDA spacetime, even for extreme parameter values where deviations from Schwarzschild geometry are significant.
@@ IV.1 Images of isoradial curves of accretion disks
As a first application (and an additional validation) we examine how the analytical formulas derived above can be used to reconstruct images of equatorial isoradial curves, i.e., curves lying in the plane $\theta=\pi/2$. These isoradial curves can be regarded as belonging to a thin accretion disk lying on the equatorial plane. In this context, the $\alpha$$\psi$ relations provide a fully analytical mapping between the emission radius and the observed direction for direct rays, enabling efficient image reconstruction without the need for full numerical ray tracing.
@@
Figure 5: The geometry of the different coordinate frames is illustrated. The orbit of fluid element $P$ at radius $r=R$ (board), the plane containing the geodesic trajectory (blue), and the plane of the distant observer with its corresponding projection (gray) are represented. Vectors used for the study of polarization in Sec. IV.2 are also shown: the fluid velocity (purple) and the equatorial projection of the magnetic field (red).
@@
With these relations, $b$ can be written as $b=b(R,\phi)$. Since our goal is to establish a direct mapping between $(R,\phi)$ and $(b,\varphi)$—or, equivalently, between $P$ and $P^{\prime}$—it is sufficient to determine a relation linking $\varphi$ and $\phi$. This relation is given by (for a concise derivation, see Appendix A of Ref. Claros and Gallo (2024)):
@@
To test the accuracy and robustness of our analytical approximations, Fig. 6 displays images of isoradial curves ($R=\text{constant}$) computed for the JP metric under different parameter configurations and viewing inclinations. Two representative sets of non-zero deformation parameters are considered: $(\epsilon_{3},\epsilon_{4})=(15,20)$ (top row) and $(\epsilon_{3},\epsilon_{4})=(10,-2)$ (bottom row). Each configuration is evaluated for three observer inclination angles, $\theta_{o}=60^{\circ}$, $70^{\circ}$, and $85^{\circ}$. The curves obtained from exact numerical integration (solid lines) are compared with the analytical approximations (23) (Approx. T, dotted line) and (24) (Approx. L, dashed line). The same set of radii used in Sec. III is adopted here.
@@
As the figure shows, the Approx. T curves are in excellent agreement with the numerically integrated ones across all configurations. This high level of consistency persists even at large viewing inclinations—up to $\theta_{o}=85^{\circ}$—and for outer radii as large as $R=4.5r_{c}$. The resulting isoradial shapes are virtually indistinguishable, with deviations well below the graphical resolution. This confirms that the full approximation (23) reliably captures both the geometric deformation induced by the inclination and the metric corrections introduced by non-zero $\epsilon_{i}$ parameters. The Approx. L curves, while still providing a good qualitative description, begin to show slight discrepancies compared to the exact and Approx. T results, particularly for the largest radii and high-inclination cases. These deviations manifest mainly as small radial displacements and minor distortions on the far side of the accretion disk, where the light rays must pass close to the black hole before reaching the observer, making light-bending effects more pronounced. Nevertheless, Approx. L remains sufficiently accurate for moderate inclinations ($\theta_{o}\lesssim 70^{\circ}$) or when only coarse image reconstruction is required.
@@
For the RZ metric, the corresponding isoradial curves are presented in Fig. 7. The same values of $R$ used in the previous section are adopted here for direct comparison. The figure illustrates how the image morphology evolves with increasing disk inclination and with variations in the metric parameters $(a_{1},a_{2},b_{1},b_{2})$. As the inclination angle $\theta_{o}$ increases, the Approx. L (24) begins to lose accuracy in reproducing the far-side region of the accretion disk, particularly for $\psi\rightarrow\pi$, where light rays travel close to the event horizon before reaching the observer. This deviation is most noticeable in the first row of Fig. 7, corresponding to the case $r_{H}=1.70$. Here, the simplified linear approximation slightly underestimates the bending of light near the black hole horizon, producing small shape distortions along the far-side contours of the disk image. In contrast, for the outer isoradial curves, the linear approximation tends to overestimate the deflection, leading to marginally larger apparent displacements in the image plane. In the case of $r_{H}=2.0$ and the given choice of parameters $a_{1},a_{2},b_{1},b_{2}$, with inclinations such that $\theta_{o}<85^{\circ}$, the approximation of Eq. (24) is indistinguishable from that obtained numerically up to radial coordinate values $r$ less than $1.5r_{c}$. By contrast, the Approx. T (23) retains a high degree of accuracy throughout the entire range of $R$ and $\theta_{o}$ values explored. Even for the extreme inclination $\theta_{o}=85^{\circ}$ and the largest emission radii $r\geq 3r_{c}$, the analytical curves remain virtually indistinguishable from the numerical reference, except for a very small deviation at $\psi\rightarrow\pi$. Therefore, the inclusion of higher-order terms in the approximation not only improves the local accuracy of the $\alpha$$\psi$ relation but also ensures excellent global consistency in the image reconstruction, capturing with a very accuracy the characteristic lensing asymmetry between the near and far sides of the accretion disk.
@@
Finally, in the case of the particular EMDA-metric, images of the isoradial curves are shown for two arbitrary sets of parameters: the first corresponding to $b_{o}=0.2$ and $\mu=0.8$, and the second to $b_{o}=0.5$ and $\mu=0.5$, for both sets, $M=1$. The curves are plotted for the same values of $R$ as in the previous cases. It is observed that as the disk inclination increases, the Approx. L (24) loses accuracy. This discrepancy is most evident in the plots with parameters $b_{o}=0.5$ and $\mu=0.5$ , particularly for high inclination angles such as $\theta_{o}=85^{\circ}$, as shown in bottom row of Fig. 8. In contrast, Approx. T (23) maintains high accuracy for all values of $R$ and for all inclination angles analyzed. It is, in fact, practically indistinguishable from the numerically calculated “exact” curve, except at an angle of $\theta_{o}=85^{\circ}$ and for a value very near the event horizon $r=1.1r_{H}$, where a slight difference can be observed.
@@
Figure 6: This figure compares isoradial curves ($R=\text{constant}$) for the Johannsen-Psaltis metric. Two sets of non-zero parameters are shown: $\epsilon_{3}=15.0,\epsilon_{4}=20.0$ (top row); $\epsilon_{3}=10.0,\epsilon_{4}=-2.0$ (bottom row). Each set is plotted for three observer inclinations: $\theta_{o}=60^{\circ}$, $70^{\circ}$ and $85^{\circ}$. The exact curves, calculated numerically (solid line) are compared with two analytical approximations: Approx. T (dotted line) and Approx. L (dashed line).
@@ IV.2 $QU$-diagrams of orbiting hotspots
We next consider a light ray confined to the $\overline{POO^{\prime}}$ plane, shown as the blue plane in the inset of Fig. 5. This plane is described by the coordinate system $(x^{\prime},y^{\prime},z^{\prime})$, where the $x^{\prime}$-axis is oriented along $\overline{OP}$, and the observer $O^{\prime}$ lies in the $x^{\prime}z^{\prime}$-plane. These spatial coordinates are complemented by a temporal coordinate $t$. The set $(t,x^{\prime},y^{\prime},z^{\prime})$, together with its associated orthonormal tetrad $\{{\bf\it\hat{t}},{\bf\it\hat{x}^{\prime}},{\bf\it\hat{y}^{\prime}},{\bf\it\hat{z}^{\prime}}\}$ defined at point $P$, specifies a locally static $G$-frame. In the $G$-frame, the photon’s four-momentum ${\bf\it k}_{(G)}$ forms an emission angle $\alpha$ with respect to the $^{\bf\it\hat{x}^{\prime}}$-axis. Since ${\bf\it k}_{(G)}$ lies entirely within the $x^{\prime}z^{\prime}$-plane, its projection along the $^{\bf\it\hat{y}^{\prime}}$-axis vanishes Claros and Gallo (2024). The relation between the $P$- and $G$-frames is given by a spatial rotation about the $^{\bf\it\hat{x}}$-axis through an angle $\xi$ (see Eq. (5.29) of Ref. Claros and Gallo (2024)). .
@@
While this polarization model can be readily applied to the metric families in Sec. III.1, this work will focus exclusively on the study of $QU$-diagrams for the particular metric (III.2.1). Three arbitrary values for the dilaton parameters $b_{0}$ are considered: $0.20,0.50$, and $0.80$. For a fixed mass, the parameter $\mu$ is uniquely determined. The corresponding ISCOs are located at $5.376464654887766$, $4.318473046963146$, and $2.946940368068108$ with orbital velocities $\beta$ of $0.5185867274$, $0.557506666$, and $0.6309930347$, respectively.
In all cases we adopt the astronomical convention. In the diagrams, colored circles mark $\phi=0$ (or $\phi=2\pi$), while squares indicate $\phi=\pi$. The orbital motion is assumed to be clockwise: starting at the circle, following the light solid curve to the square, and returning to the circle along the dark solid curve.
@@ V Final Remarks
Specifically, we have derived fully analytical expressions for two widely used parametric metric families—the Johannsen–Psaltis and Rezzolla–Zhidenko metrics—allowing a direct mapping between emission and observation points for asymptotic observers. These analytical expressions make it possible to perform fast and accurate ray-tracing computations across different gravity models. Furthermore, we have obtained compact analytical formulas for the Einstein–Maxwell–dilaton–axion spacetime and applied them to study synchrotron polarization through synthetic $QU$-loop diagrams.

### v2

@@ III.2.1 Einstein-Maxwell-dilaton-axion metric
Figure 4 illustrates the performance of the analytical approximations given by Eqs. (23) and (24) for the Einstein–Maxwell–dilaton–axion metric, considering three representative configurations corresponding to different combinations of the parameters $(b_{0},\mu)$ and the associated $(r_{H},r_{c})$ values. The plots display the relation between $\cos\psi$ and $\cos\alpha$ for several emission radii $R$, together with the relative percentage deviation of $\alpha$ in the lower panels. Across all parameter choices, the full analytical approximation given by Eq. (23) (Approx. T) shows excellent agreement with the exact numerical integration. The discrepancy remains within $|\delta\alpha|/\alpha<3\%$ over the entire domain of $\psi$, even for the smallest radii considered; namely, $R=1.1r_{H}$ and $R=1.01r_{c}$; where strong deflection effects dominate. This shows that the approximation maintains high accuracy even in the deep strong-field regime near the event horizon. The linearized expression derived from Eq. (24) (‘Approx. L’) also performs well for moderate deflections ($\cos\psi\gtrsim 0$), particularly at larger radii ($R\geq 1.5r_{c}$), but gradually loses precision as $\psi\to\pi$ and $\cos\psi\to-1$. This degradation is more evident in the rightmost panel, where the dilaton parameter dominates ($b_{0}>\mu$), leading to a more compact horizon and stronger curvature gradients. In contrast, in the leftmost configuration ($b_{0}<\mu$), both approximations remain nearly indistinguishable from the numerical result except for very large bending angles. As in the previous cases, the lower panels quantify these trends. For all three configurations, ‘Approx. T’ maintains a smooth and nearly symmetric error distribution around $\cos\psi=0$, indicating robustness under parameter variations. ‘Approx. L’, while showing larger residuals near $\cos\psi\approx-1$, still provides a reasonable description in the weak-to-intermediate deflection regime, making it suitable for analytical estimates where computational simplicity is preferred. Overall, these results show that the generalized analytical formulas accurately reproduces light propagation in the EMDA spacetime, even for extreme parameter values where deviations from Schwarzschild geometry are significant.
@@ IV.1 Images of isoradial curves of accretion disks
As a first application (and an additional validation) we examine how the analytical formulas derived above can be used to reconstruct images of equatorial isoradial curves, i.e., curves of fixed $r$ lying in the plane $\theta=\pi/2$. These isoradial curves can be regarded as belonging to a thin accretion disk lying on the equatorial plane. In this context, the $\alpha$$\psi$ relations provide a fully analytical mapping between the emission radius and the observed direction for direct rays, enabling efficient image reconstruction without the need for full numerical ray tracing.
@@
Figure 5: Figure reproduced from Ref. Claros and Gallo (2024), Fig. 7. The geometry of the different coordinate frames is illustrated. The orbit of fluid element $P$ at radius $r=R$ (board), the plane containing the geodesic trajectory (blue), and the plane of the distant observer with its corresponding projection (gray) are represented. Vectors used for the study of polarization in Sec. IV.2 are also shown: the fluid velocity (purple) and the equatorial projection of the magnetic field (red).
@@
With these relations, $b$ can be written as $b=b(R,\phi)$. Since our goal is to establish a direct mapping between $(R,\phi)$ and $(b,\varphi)$-or, equivalently, between $P$ and $P^{\prime}$-it is sufficient to determine a relation linking $\varphi$ and $\phi$. This relation is given by (for a concise derivation, see Appendix A of Ref. Claros and Gallo (2024)):
@@
To test the accuracy and robustness of our analytical approximations, Fig. 6 displays images of isoradial curves ($R=\text{constant}$) computed for the JP metric under different parameter configurations and viewing inclinations. Two representative sets of non-zero deformation parameters are considered: $(\epsilon_{3},\epsilon_{4})=(15,20)$ (top row) and $(\epsilon_{3},\epsilon_{4})=(10,-2)$ (bottom row). Each configuration is evaluated for three observer inclination angles, $\theta_{o}=60^{\circ}$, $70^{\circ}$, and $85^{\circ}$. The curves obtained from exact numerical integration (solid lines) are compared with the analytical approximations (23) (‘Approx. T’, dotted line) and (24) (‘Approx. L’, dashed line). The same set of radii used in Sec. III is adopted here.
@@
As the figure shows, the ‘Approx. T’ curves are in excellent agreement with the numerically integrated ones across all configurations. This high level of consistency persists even at large viewing inclinations; up to $\theta_{o}=85^{\circ}$; and for outer radii as large as $R=4.5r_{c}$. The resulting isoradial shapes are virtually indistinguishable, with deviations well below the graphical resolution. This confirms that the full approximation (23) reliably captures both the geometric deformation induced by the inclination and the metric corrections introduced by non-zero $\epsilon_{i}$ parameters. The ‘Approx. L’ curves, while still providing a good qualitative description, begin to show slight discrepancies compared to the exact and ‘Approx. T’ results, particularly for the largest radii and high-inclination cases. These deviations manifest mainly as small radial displacements and minor distortions on the far side of the accretion disk, where the light rays must pass close to the black hole before reaching the observer, making light-bending effects more pronounced. Nevertheless, ‘Approx. L’ remains sufficiently accurate for moderate inclinations ($\theta_{o}\lesssim 70^{\circ}$) or when only coarse image reconstruction is required.
@@
For the RZ metric, the corresponding isoradial curves are presented in Fig. 7. The same values of $R$ used in the previous section are adopted here for direct comparison. The figure illustrates how the image morphology evolves with increasing disk inclination and with variations in the metric parameters $(a_{1},a_{2},b_{1},b_{2})$. As the inclination angle $\theta_{o}$ increases, the ‘Approx. L’ (24) begins to lose accuracy in reproducing the far-side region of the accretion disk, particularly for $\psi\rightarrow\pi$, where light rays travel close to the event horizon before reaching the observer. This deviation is most noticeable in the first row of Fig. 7, corresponding to the case $r_{H}=1.70$. Here, the simplified linear approximation slightly underestimates the bending of light near the black hole horizon, producing small shape distortions along the far-side contours of the disk image. In contrast, for the outer isoradial curves, the linear approximation tends to overestimate the deflection, leading to marginally larger apparent displacements in the image plane. In the case of $r_{H}=2.0$ and the given choice of parameters $a_{1},a_{2},b_{1},b_{2}$, with inclinations such that $\theta_{o}<85^{\circ}$, the approximation of Eq. (24) is indistinguishable from that obtained numerically up to radial coordinate values $r$ less than $1.5r_{c}$. By contrast, the ‘Approx. T’ (23) retains a high degree of accuracy throughout the entire range of $R$ and $\theta_{o}$ values explored. Even for the extreme inclination $\theta_{o}=85^{\circ}$ and the largest emission radii $r\geq 3r_{c}$, the analytical curves remain virtually indistinguishable from the numerical reference, except for a very small deviation at $\psi\rightarrow\pi$. Therefore, the inclusion of higher-order terms in the approximation not only improves the local accuracy of the $\alpha$$\psi$ relation but also ensures excellent global consistency in the image reconstruction, capturing with a very accuracy the characteristic lensing asymmetry between the near and far sides of the accretion disk.
@@
Finally, in the case of the particular EMDA-metric, images of the isoradial curves are shown for two arbitrary sets of parameters: the first corresponding to $b_{o}=0.2$ and $\mu=0.8$, and the second to $b_{o}=0.5$ and $\mu=0.5$, for both sets, $M=1$. The curves are plotted for the same values of $R$ as in the previous cases. It is observed that as the disk inclination increases, the ‘Approx. L’ (24) loses accuracy. This discrepancy is most evident in the plots with parameters $b_{o}=0.5$ and $\mu=0.5$ , particularly for high inclination angles such as $\theta_{o}=85^{\circ}$, as shown in bottom row of Fig. 8. In contrast, ‘Approx. T’ (23) maintains high accuracy for all values of $R$ and for all inclination angles analyzed. It is, in fact, practically indistinguishable from the numerically calculated “exact” curve, except at an angle of $\theta_{o}=85^{\circ}$ and for a value very near the event horizon $r=1.1r_{H}$, where a slight difference can be observed.
@@
Figure 6: This figure compares isoradial curves ($R=\text{constant}$) for the Johannsen-Psaltis metric. Two sets of non-zero parameters are shown: $\epsilon_{3}=15.0,\epsilon_{4}=20.0$ (top row); $\epsilon_{3}=10.0,\epsilon_{4}=-2.0$ (bottom row). Each set is plotted for three observer inclinations: $\theta_{o}=60^{\circ}$, $70^{\circ}$ and $85^{\circ}$. The exact curves, calculated numerically (solid line) are compared with two analytical approximations: ‘Approx. T’ (dotted line) and ‘Approx. L’ (dashed line).
@@ IV.2 $QU$-diagrams of orbiting hotspots
We next consider a light ray confined to the $\overline{POO^{\prime}}$ plane, shown as the blue plane in the inset of Fig. 5. This plane is described by the coordinate system $(x^{\prime},y^{\prime},z^{\prime})$, where the $x^{\prime}$-axis is oriented along $\overline{OP}$, and the observer $O^{\prime}$ lies in the $x^{\prime}z^{\prime}$-plane. These spatial coordinates are complemented by a temporal coordinate $t$. The set $(t,x^{\prime},y^{\prime},z^{\prime})$, together with its associated orthonormal tetrad $\{{\bf\it\hat{t}},{\bf\it\hat{x}^{\prime}},{\bf\it\hat{y}^{\prime}},{\bf\it\hat{z}^{\prime}}\}$ defined at point $P$, specifies a locally static $G$-frame. In the $G$-frame, the photon’s four-momentum ${\bf\it k}_{(G)}$ forms an emission angle $\alpha$ with respect to the $^{\bf\it\hat{x}^{\prime}}$-axis. Since ${\bf\it k}_{(G)}$ lies entirely within the $x^{\prime}z^{\prime}$-plane, its projection along the $^{\bf\it\hat{y}^{\prime}}$-axis vanishes Claros and Gallo (2024). The relation between the $P$- and $G$-frames is given by a spatial rotation about the $^{\bf\it\hat{x}}$-axis through an angle $\xi$ (see Eq. (5.29) of Ref. Claros and Gallo (2024)).
@@
While this polarization model can be readily applied to the metric families in Sec. III.1, this work will focus exclusively on the study of $QU$-diagrams for the particular metric (III.2.1). Three representative values of the dilaton parameter $b_{0}$ are considered: $0.20$, $0.50$, and $0.80$. For a fixed mass, the parameter $\mu$ is uniquely determined. The corresponding ISCO radii are located approximately at $5.376$, $4.318$, and $2.946$, with the corresponding orbital velocities $\beta$ being approximately $0.518$, $0.557$, and $0.631$, respectively.
In all cases, we adopt the standard radio-astronomical convention: the polarization angle is measured from the $Y^{\prime}$-axis in a counterclockwise direction with $Q$ and $U$ defined as minus the expressions in Eqs. (45) and (46) (see, for example, Fig. 1 of Vincent et al. (2024)). In the diagrams, colored circles mark $\phi=0$ (or $\phi=2\pi$), while squares indicate $\phi=\pi$. The orbital motion is assumed to be clockwise: starting at the circle, following the light solid curve to the square, and returning to the circle along the dark solid curve.
@@ V Final Remarks
Specifically, we have derived fully analytical expressions for two widely used parametric metric families-the Johannsen–Psaltis and Rezzolla–Zhidenko metrics-allowing a direct mapping between emission and observation points for asymptotic observers. These analytical expressions make it possible to perform fast and accurate ray-tracing computations across different gravity models. Furthermore, we have obtained compact analytical formulas for the Einstein–Maxwell–dilaton–axion spacetime and applied them to study synchrotron polarization through synthetic $QU$-loop diagrams.